OpenStudy (anonymous):

2|3(x - 5)| + 3 = 9 (NOT A TRICK QUESTION)

OpenStudy (anonymous):

x=4.6

OpenStudy (angela210793):

2|3(x - 5)| + 3 = 9 - 3x+15=3 2|3(x - 5)| = 9 - 3=6 -3x=-12 |3(x - 5)|=6/2=3 x=4 3x-15=3 3x=18 x=6

OpenStudy (anonymous):

i.e my reply is 4,6 not 4.6 sry .. i'm so sleepy right now sorry :)